One of the tallest buildings in the world is the Taipei 101 in Taiwan, at a height of 1671 feet. Assume that this height was measured on a cool spring day when the temperature was 15.50C. You could use the building as a sort of giant thermometer on a hot summer day by carefully measuring its height. Suppose you do this and discover that the Taipei 101 is 0.471 foot taller than its official height. What is the temperature, assuming that the building is in thermal equilibrium with the air and that its entire frame is made of steel? (coefficient of linear expansion of steel is 3.5x10-5/0C)


1
Expert's answer
2021-10-31T18:12:51-0400

Determine the temperature.


"\Delta L=L_0\alpha(T_f-T_0),\\\space\\
T_f=\frac{\Delta L}{L_0\alpha}+T_0=23.6°\text C."